The sun rises at 5:00 AM and is directly overhead at 11:00 AM. At what time in the morning (to the nearest minute) does a 42m tall redwood tree cast a 50m shadow? A person standing next to the tree casts a 1.7m shadow at the same time that the tree is casting the 50m shadow. Find the height of the person.
Answer by stanbon(75887)

The sun rose 90 degrees in 6 hrs.
Draw a triangle with the 42m tall tree and its 50m long shadow.
Using the inverse tangent you find the elevation angle of the sun
is 40.03 degrees above the horizon.
Multiply (40.03/90)(6 hrs)= 2.668 hr= 2hr and 40.08minutes
The time the shadow was 50m long was (5 hr + 2 hr 40min)= 7:40
The person standing next to the tree is x meters high with shadow 1.7 meters.
Therefore 42/50 = x/1.7
x=1.7(42/50)=1.428 meters
Cheers,
stan H.